OpenStudy (openstudier):

Every quiz score is worth 1/20 of a student’s grade at the end of the quarter. Every student will have four quarter grades by the end of the year and each of these quarter grades is worth 2/9 of the student’s Final Year-End grade. The final exam, taken on the last day of school, will be worth 1/9 of the Final Year-End grade. According to this grading policy, the final exam has the same importance (weight) as how many quizzes?

OpenStudy (solomonzelman):

Each quarter is 20 quizzes. Each quarter is 2/9 of the early grade. Thus, Each quarter = 20 quizzes 20 quizzess = 2/9 of yearly grade If final exam is 1/9 of yearly grade, then how many quizzes is 1/9 of yearly grade equal to?

OpenStudy (solomonzelman):

Basically just using the principle \(\color{#000000 }{ \displaystyle a=b,~b=c,~\therefore~a=c }\) ( \(\color{#000000 }{ \displaystyle\therefore }\) means "therefore" ) you should be able to answer my question...

OpenStudy (solomonzelman):

That is as simple as: 2 apples are worth $10, then 1 apple is worth $10/2=$5.